# Linguistic Collapse: Neural Collapse in (Large) Language Models

### v2

@@ 6 Limitations
While to the best of our knowledge, no previous work has studied realistic stochastic token prediction, it is possible that the quantities that we measure are not perfectly suited for $\mathcal{NC}$ in language modeling. As we described in §1.1, the $\mathcal{NC}$ framework does not translate neatly to the language modeling space due to many adverse conditions, so full convergence to $\mathcal{NC}$ in the TPT was highly improbable. This paper leaves much room for future work to better adapt $\mathcal{NC}$ for next-token prediction, which we discuss further in Section 7.
Language modeling
@@ 7 Discussion
Past works have established that properties resembling $\mathcal{NC}$ evolve as a function of model depth [4, 36, 53, 59, 60, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105]. Layer-wise $\mathcal{NC}$ – sometimes dubbed deep neural collapse ($\mathcal{DNC}$) [53, 59] – and related phenomena at intermediate layers remain an interesting subtopic. We leave their study and induction in CLMs (like [106]) as future work.
@@
Given the evidence for the development of $\mathcal{NC}$ and associated generalization under various loss functions [19, 49, 60, 101, 107, 108, 109] in other domains, NLP researchers may still benefit from analyzing, adapting or training towards $\mathcal{NC}$. As alluded to earlier, the simplex ETF and even the CE loss may not be truly optimal for this problem setting, so we anticipate future works to both construct more amenable geometries with better-suited objectives and then capitalize on their benefits downstream. As discussed in §2, there is an abundance of literature in $\mathcal{NC}$, some of which could potentially adapt $\mathcal{NC}$ to be useful for NLP; we hope to inspire such investigations.
@@
Researchers in NLP and multimodal settings are ultimately interested in measuring model performance on practical tasks. Notable benchmarks include GLUE [77], MMLU [110], and BIG-bench [111]. However, several contemporaries [112, 113, 114] have demonstrated that models’ downstream capabilities are roughly correlated with their abilities to effectively compress their pre-training data. Based on their findings, our application of the $\mathcal{NC}$ framework to the pre-training stage of CLMs against validation CE loss should be an appropriate first step in this intersection. Looking forward, we anticipate exciting analysis for language modeling tasks or benchmarks, especially on creativity and retrieval for natural language understanding and generation.
Conversely, some form of $\mathcal{NC}$ could even serve as an alternative evaluation. Although it would be prohibitively expensive to measure $\mathcal{NC}$ on the vast and sometimes obscure pre-training data of most frontier production LLMs, doing so on a small set of unseen data (i.e. test data) would be realistic.
@@
At a high level, the number and density of clusters for a token can reflect its learned meanings and uses. This would be particularly useful as LLMs adapt to ever-evolving language and further expansion into non-English domains. Our formulae (Section 3) and results (Section 4) expose the pairwise token class interactions in noise, interference, classifier duality, and classifier agreement in the top-level features. Similarly to works in other domains [21, 55, 72, 115], these $\mathcal{NC}$ metrics can serve as a form of low-level interpretability to aid understanding certain behaviors of (L)LMs. Between tokens, one can often discern how related or interchangeable words are based on their pair-wise interactions, or how antithetical or unrelated they are based on orthogonality. For example, we present a rudimentary analysis of homonyms and English first names in Appendix Q.
@@ 8 Conclusion
In the short term, our work presents rudimentary techniques to analyze and interpret token-level properties of (L)LMs. We anticipate future work to suitably adapt $\mathcal{NC}$ (and related frameworks) to the still-fresh frontier of autoregressive language modeling. Researchers could then effectively capitalize on previous learnings from $\mathcal{NC}$ to better understand and improve the pre/post-training processes of increasingly complex and large language (multimodal) models.
@@ Acknowledgements
We thank Elliot Creager, David Glukhov, Daniel Johnson, Jivan Waber, and Colin Raffel for their helpful feedback and stimulating discussions. Aditya Mehrotra and Yu Bo Gao provided technical assistance in our implementations. We acknowledge the support of Canada’s Natural Sciences and Engineering Research Council (NSERC, www.nserc-crsng.gc.ca/). This research was enabled in part by resources from Calcul Québec (www.calculquebec.ca), the Digital Research Alliance of Canada (www.alliancecan.ca), and the Vector Institute (www.vectorinstitute.ai).
@@ Appendix A Dataset
TinyStories is a dataset of short children’s stories generated by GPT-3.5 and GPT-4 [2], released with the CDLA-Sharing-1.0 licence. We trained and evaluated on the first version, as described:
@@ A.1 Alternative (Real) Datasets
The study of $\mathcal{NC}$ in causal language modeling at the token level would be very expensive, so the motivation to use a small dataset is clear. However, most commonly used text datasets such as WikiText [117], BookCorpus [118], CommonCrawl [119], or most subsets from the Pile [120] are much too complex and broad to be effectively compressed by CLMs of the scale that we work with.
@@ A.2 On the Use of TinyStories
According to its authors, TinyStories [2] is explicitly designed to preserve the essential elements of natural language, such as grammar, vocabulary, facts, and reasoning, while being smaller and more refined in terms of its breadth and diversity. Unlike large corpora that can overwhelm small language models (SLMs) due to their excessive breadth and diversity, TinyStories offers a concentrated dataset that hones in on core linguistic structures and reasoning capabilities. This is evident in its small word vocabulary, consisting of approximately 1500 words that a child would use, and in its 29K empirical vocabulary under the GPT-Neo tokenizer.
Despite its concentrated nature, TinyStories enables models trained on it to produce grammatically correct, factual, and reasonable stories. Additionally, these models can be finetuned on specific instructions found in the TinyStories-Instruct dataset. The authors of TinyStories also demonstrate that their models can creatively produce stories that are dissimilar enough to their training data, indicating a balanced capability for generalization and creativity.
@@
Furthermore, TinyStories was created using GPT 3.5/4, which are advanced language models with significantly larger architectures trained on orders of magnitude more tokens, helping minimize the effect of the synthetic nature of the generated dataset. We also considered a possible effect of model collapse as a result of training on synthetic data [121] and follow-up work [122] suggest that a single iteration of data generation (as generated TinyStories) has a very negligible model collapse.
@@ Appendix C Optimization
Training was performed using an adaptation of an open-source causal language modeling script from Huggingface: https://github.com/huggingface/transformers/blob/main/examples/pytorch/language-modeling/run_clm.py
@@ Appendix D Embeddings Collection & $\mathcal{NC}$ Analysis
One pass over the train set for embeddings collection can take up to 6 hours on a single NVIDIA A100 (40GB) GPU. Analysis of a single metric for a given model takes no longer than 5 minutes.
Appendix E Within-Class Variability Collapse with Scale – $\mathcal{NC}_{1}$
@@
Appendix F Mean Norms Growth with Scale – (Related to $\mathcal{NC}_{2}$)
Figure 8: Class embedding mean norms grow when scaling width ($d$) and across training for 1 (left) through 10 (right) epochs with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom).
Figure 9: Class embedding mean norms grow when scaling depth ($L$) and across training for 1 (left) through 10 (right) epochs with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom).
Appendix G Equinormness with Scale – $\mathcal{(G)NC}_{2}$
Figure 10: Variation in embedding norms decreases ($\mathcal{NC}_{2}$) when scaling width ($d$) in models trained for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom). Note that the degree of equinormness eventually plateaus for sufficiently deep and trained models.
Figure 11: Variation in embedding norms decreases ($\mathcal{NC}_{2}$) when scaling depth ($L$) in models trained for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom).
Appendix H Interference with Scale – (Related to $\mathcal{NC}_{2}$)
@@
Appendix I Equiangularity with Scale – (Against $\mathcal{NC}_{2}$)
@@
Appendix J Logarithmic Distances with Scale – (Related to $\mathcal{GNC}_{2}$)
@@
Appendix K Hyperspherical Uniformity with Scale – $\mathcal{GNC}_{2}$
@@
Figure 20: Generalization (validation loss) shows some correlation with average mean norms (left) and almost none with their variations (i.e. equinormness, $\mathcal{NC}_{2}$) (right).
@@
Appendix M Self-Duality with Scale – (Against $\mathcal{NC}_{3}$)
@@
Figure 23: Average classifier alignment increases $\mathcal{NC}$3 when training for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom). However, we see no meaningful trend when scaling width $d$, suggesting that $\mathcal{NC}$3 does not coalesce with language modeling training.
Figure 24: Average classifier alignment increases $\mathcal{NC}$3 when training for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom). However, we see no meaningful trend when scaling depth $L$, suggesting that $\mathcal{NC}$3 does not coalesce with language modeling training.
Appendix N Uniformity Duality with Scale – $\mathcal{UNC}_{3}$
@@
Appendix P Classifier Agreement – $\mathcal{NC}_{4}$
@@
This section presents token-wise interpretability results found in top-layer embeddings from our most learned model (https://huggingface.co/rhubarbwu/TinyStories-12x1024_10L).

### v3

@@ 6 Limitations
While to the best of our knowledge, no previous work has studied realistic stochastic token prediction, it is possible that the quantities that we measure are not perfectly suited for $\mathcal{NC}$ in language modelling. As we described in §1.1, the $\mathcal{NC}$ framework does not translate neatly to the language modelling space due to many adverse conditions, so full convergence to $\mathcal{NC}$ in the TPT was highly improbable. This paper leaves much room for future work to better adapt $\mathcal{NC}$ for next-token prediction, which we discuss further in Section 7.
Language modelling
@@ 7 Discussion
Past works have established that properties resembling $\mathcal{NC}$ evolve as a function of model depth [4, 36, 53, 56, 59, 60, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104]. Layer-wise $\mathcal{NC}$ — sometimes dubbed deep neural collapse ($\mathcal{DNC}$) [53, 59] — and related phenomena at intermediate layers remain an interesting subtopic. We leave their study and induction in CLMs (like [105]) as future work.
@@
Given the evidence for the development of $\mathcal{NC}$ and associated generalization under various loss functions [19, 49, 60, 100, 106, 107, 108] in other domains, NLP researchers may still benefit from analyzing, adapting or training towards $\mathcal{NC}$. As alluded to earlier, the simplex ETF and even the CE loss may not be truly optimal for this problem setting, so we anticipate future works to both construct more amenable geometries with better-suited objectives and then capitalize on their benefits downstream. As discussed in §2, there is an abundance of literature in $\mathcal{NC}$, some of which could potentially adapt $\mathcal{NC}$ to be useful for NLP; we hope to inspire more such investigations.
@@
Researchers in NLP and multimodal settings are ultimately interested in measuring model performance on practical tasks; notable benchmarks include GLUE [77], MMLU [109], and BIG-bench [110]. However, several contemporaries [111, 112, 113] have demonstrated that models’ downstream capabilities are roughly correlated with their abilities to effectively compress their pre-training data. Based on their findings, our application of the $\mathcal{NC}$ framework to the pre-training stage of CLMs against validation CE loss should be an appropriate first step in this intersection. Looking forward, we anticipate exciting analysis for language modelling tasks or benchmarks, especially on creativity and retrieval for natural language understanding and generation.
Conversely, some form of $\mathcal{NC}$ could be an alternative evaluation. Although it would be prohibitively expensive to measure $\mathcal{NC}$ on the vast and sometimes obscure pre-training data of most frontier production LLMs, doing so on a small set of in-distribution data (i.e. test set) would be realistic.
@@
At a high level, the number and density of clusters for a token can reflect its learned meanings and uses. This would be particularly useful as LLMs adapt to ever-evolving language and further expansion into non-English domains. Our formulae (Section 3) and results (Section 4) expose the pairwise token class interactions in noise, interference, classifier duality, and classifier agreement in the top-level features. Similarly to works in other domains [21, 55, 72, 114], these $\mathcal{NC}$ metrics can serve as a form of low-level interpretability to aid understanding certain behaviours of (L)LMs. Between tokens, one can often discern how related or interchangeable words are based on their pair-wise interactions, or how antithetical or unrelated they are based on orthogonality. For example, we present a rudimentary analysis of homonyms and English first names in Appendix Q.
@@ 8 Conclusion
In the short term, our work presents rudimentary techniques to analyze and interpret token-level properties of (L)LMs. We anticipate future work to suitably adapt $\mathcal{NC}$ (and related frameworks) to the still-fresh frontier of autoregressive language modelling. Researchers could then effectively capitalize on previous learnings from $\mathcal{NC}$ to better understand and improve the pre/post-training processes of increasingly complex and large language (multimodal) models.
@@ Acknowledgements
We thank Elliot Creager, David Glukhov, Daniel Johnson, Jivan Waber, and Colin Raffel for their helpful feedback and stimulating discussions. Aditya Mehrotra and Yu Bo Gao provided technical assistance in our implementations. We acknowledge the support of the Natural Sciences and Engineering Research Council (NSERC) of Canada (www.nserc-crsng.gc.ca/). This research was enabled in part by resources from Calcul Québec (www.calculquebec.ca), the Digital Research Alliance of Canada (www.alliancecan.ca), and the Vector Institute (www.vectorinstitute.ai).
@@ Appendix A Dataset
TinyStories is a dataset of short children’s stories generated by GPT-3.5 and GPT-4 [2], released with the CDLA-Sharing-1.0 licence. We trained and evaluated models on their first version, as described:
@@ A.1 Alternative (Real) Datasets
The study of $\mathcal{NC}$ in causal language modelling at the token level would be unreasonably expensive, so the motivation to use a small dataset is clear. However, most commonly used text datasets such as WikiText [116], BookCorpus [117], CommonCrawl [118], or most subsets from the Pile [119] are much too complex and broad to be effectively compressed by CLMs of the scale that we work with.
@@ A.2 On the Use of TinyStories
According to its authors, TinyStories [2] is explicitly designed to preserve the essential elements of natural language, such as grammar, vocabulary, facts, and reasoning, while being smaller and more refined in terms of its breadth and diversity. Unlike large corpora that can overwhelm small language models (SLMs) due to their excessive breadth and diversity, TinyStories offers a concentrated dataset that hones in on core linguistic structures and reasoning capabilities. This is evident in its small vocabulary, consisting of approximately 1500 words that a child would use, and in its 29K empirical vocabulary under the GPT-Neo tokenizer.
Despite its concentrated nature, TinyStories enables models trained on it to produce grammatically correct, factual, and reasonable stories. Additionally, these models can be finetuned on specific instructions found in the TinyStories-Instruct dataset. The authors of TinyStories also demonstrate that their models can creatively produce stories dissimilar enough to their training data, indicating a balanced capability for generalization and creativity.
@@
Furthermore, TinyStories was created using GPT 3.5/4, advanced language models with significantly larger architectures trained on orders of magnitude more tokens; this should help minimize the effect of the synthetic nature of the generated dataset. We also considered a possible effect of model collapse as a result of training on synthetic data [120] and follow-up work [121] suggest that a single iteration of data generation (as generated TinyStories) has a very negligible model collapse.
@@ Appendix C Optimization
The training was performed using an adaptation of an open-source causal language modelling script from Huggingface: https://github.com/huggingface/transformers/blob/main/examples/pytorch/language-modeling/run_clm.py
@@ Appendix D Embeddings Collection & $\mathcal{NC}$ Analysis
One pass over the train set for embeddings collection can take up to 6 hours on a single NVIDIA A100 (40GB) GPU. Analysis of a single metric for a given model takes less than 5 minutes.
Appendix E Within-Class Variability Collapse with Scale — $\mathcal{NC}_{1}$
@@
Appendix F Mean Norms Growth with Scale — (Related to $\mathcal{NC}_{2}$)
Figure 8: Logarithmic class mean norms grow when scaling width ($d$) and across training for 1 (left) through 10 (right) epochs with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom).
Figure 9: Logarithmic class mean norms grow when scaling depth ($L$) and across training for 1 (left) through 10 (right) epochs with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom).
Appendix G Equinormness with Scale — $\mathcal{(G)NC}_{2}$
Figure 10: Variation in (logarithmic) norms decreases ($\mathcal{NC}_{2}$) when scaling width ($d$) in models trained for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom). Note that the degree of equinormness eventually plateaus for sufficiently deep and trained models.
Figure 11: Variation in (logarithmic) norms decreases ($\mathcal{NC}_{2}$) when scaling depth ($L$) in models trained for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom).
Appendix H Interference with Scale — (Related to $\mathcal{NC}_{2}$)
@@
Appendix I Equiangularity with Scale — (Against $\mathcal{NC}_{2}$)
@@
Appendix J Logarithmic Distances with Scale — (Related to $\mathcal{GNC}_{2}$)
@@
Appendix K Hyperspherical Uniformity with Scale — $\mathcal{GNC}_{2}$
@@
Figure 20: Generalization (validation loss) shows some correlation with logarithmic mean norms in both their average (left) and variations (i.e. equinormness, $\mathcal{NC}_{2}$) (right).
@@
Appendix M Self-Duality with Scale — (Against $\mathcal{NC}_{3}$)
@@
Figure 23: Average classifier alignment increases $\mathcal{NC}$3 when training for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom). However, we see no meaningful trend when scaling width $d$, suggesting that $\mathcal{NC}$3 does not coalesce with language modelling training.
Figure 24: Average classifier alignment increases $\mathcal{NC}$3 when training for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom). However, we see no meaningful trend when scaling depth $L$, suggesting that $\mathcal{NC}$3 does not coalesce with language modelling training.
Appendix N Uniformity Duality with Scale — $\mathcal{UNC}_{3}$
@@
Appendix P Classifier Agreement — $\mathcal{NC}_{4}$
@@
This section presents token-wise interpretability results from top-layer embeddings from our most learned model (https://huggingface.co/rhubarbwu/TinyStories-12x1024_10L).